# Program evaluation with remotely sensed outcomes

### v2

@@ Appendix C Identification proofs
C.2 Proof of Lemma 3.1
Let $\delta_{y,d}^{e}(r,x):=\Pr(R=r\mid Y=y,D=d,S=e,X=x)$. By the law of total probability,
$\delta_{d}^{e}(r,x)$ $=\delta_{0,d}^{e}(r,x)+\left\{\delta_{1,d}^{e}(r,x)-\delta_{0,d}^{e}(r,x)\right\}\Pr\{Y(d)=1\mid D=d,S=e,X=x\}$
$=\delta_{0,d}^{e}(r,x)+\left\{\delta_{1,d}^{e}(r,x)-\delta_{0,d}^{e}(r,x)\right\}\Pr\{Y(d)=1\mid S=e,X=x\}$
where the second equality applies Assumption 2.1. Next, notice that
$\Pr(R=r\mid Y=y,D=d,S=e,X=x)$ $=\Pr(R=r\mid Y=y,D=d,S=o,X=x)$
$=\Pr(R=r\mid Y=y,S=o,X=x),$
where the first equality applies Assumption 2.2 and the second equality applies Assumption 2.3(ii). Combining the previous two displays,
$\delta_{d}^{e}(r,x)=\delta_{0}^{o}(r,x)+\left\{\delta_{1}^{o}(r,x)-\delta_{0}^{o}(r,x)\right\}\Pr\{Y(d)=1\mid S=e,X=x\}.\qed$
C.3 Proof of Theorem 3.1
To prove this result, we apply Bayes’ rule to rewrite
$\delta_{y}^{o}(r,x)$ $:=\Pr(R=r\mid Y=y,S=o,X=x)=\frac{\Pr(Y=y,S=o\mid R=r,X=x)\Pr(R=r|X=x)}{\Pr(Y=y,S=o|X=x)},$
$\delta_{d}^{e}(r,x)$ $:=\Pr\left(R=r\mid D=d,S=e,X=x\right)=\frac{\Pr(D=d,S=e\mid R=r,X=x)\Pr(R=r|X=x)}{\Pr(D=d,S=e|X=x)}.$
Applying Lemma 3.1, we have for $d\in\{0,1\},x\in\mathcal{X}$ and $r\in\mathcal{R}$,
$\frac{\Pr(D=d,S=e\mid R=r,X=x)}{\Pr(D=d,S=e|X=x)}-\frac{\Pr(Y=0,S=o\mid R=r,X=x)}{\Pr(Y=0,S=o|X=x)}=$
$\left\{\frac{\Pr(Y=1,S=o\mid R=r,X=x)}{\Pr(Y=1,S=o|X=x)}-\frac{\Pr(Y=0,S=o\mid R=r,X=x)}{\Pr(Y=0,S=o|X=x)}\right\}\mu(d,x).$
@@
$\mathbb{E}\left[\frac{1\left\{D=d,S=e\right\}}{\Pr(D=d,S=e|X=x)}-\frac{1\left\{Y=y,S=o\right\}}{\Pr(Y=y,S=o|X=x)}\mid R=r,X=x\right]=$
$\mathbb{E}\left[\frac{1\left\{Y=1,S=o\right\}}{\Pr(Y=1,S=o)}-\frac{1\left\{Y=0,S=o\right\}}{\Pr(Y=1,S=o)}\mid R=r,X=x\right]\mu(d,x).$
@@
$\mathbb{E}\left[\frac{1\left\{D=1,S=e\right\}}{\Pr(D=1,S=e)}-\frac{1\left\{D=0,S=e\right\}}{\Pr(D=0,S=e)}\mid R=r,X=x\right]=$
$\mathbb{E}\left[\frac{1\left\{Y=1,S=o\right\}}{\Pr(Y=1,S=o)}-\frac{1\left\{Y=0,S=o\right\}}{\Pr(Y=1,S=o)}\mid R=r,X=x\right]\theta(x),$
@@
C.4 Proof of Theorem 3.2
The proof follows the same steps as the proofs of Lemma 3.1 and Theorem 3.1. Applying the law of total probability, for $d\in\{0,1\}$ and $r\in\mathcal{R}$, we have
$\delta_{d}^{e}(r,x)=\delta_{0,d}^{e}(r,x)+\left\{\delta_{1,d}^{e}(r,x)-\delta_{0,d}^{e}(r,x)\right\}\Pr\{Y(d)=1\mid S=e,X=x\}$
for $\delta_{y,d}^{e}(r,x):=\Pr(R=r\mid Y=y,D=d,S=e,X=x)$. By Assumptions 2.2 and 2.3(i),
$\Pr(R=r\mid Y=y,D=d,S=e,X=x)=\Pr(R=r\mid Y=y,D=d,S=o,X=x).$
Combining the previous two displays, we have
$\delta_{d}^{e}(r,x)=\delta_{0,d}^{o}(r,x)+\left(\delta_{1,d}^{o}(r,x)-\delta_{0,d}^{o}(r,x)\right)\Pr\{Y(d)=1\mid S=e,X=x\}.$
We next apply Bayes’ rule to rewrite
$\delta_{d}^{e}(r,x)=\frac{\Pr(D=d,S=e\mid R=r,X=x)\Pr(R=r|X=x)}{\Pr(D=d,S=e|X=x)}$
$\delta_{y,d}^{o}(r,x)=\frac{\Pr(Y=y,D=d,S=o\mid R=r,X=x)\Pr(R=r|X=x)}{\Pr(Y=y,D=d,S=o|X=x)}.$
It therefore follows that, for $d\in\{0,1\}$ and $r\in\mathcal{R}$,
$\frac{\Pr(D=d,S=e\mid R=r,X=x)}{\Pr(D=d,S=e|X=x)}-\frac{\Pr(Y=0,D=d,S=o\mid R=r,X=x)}{\Pr(Y=0,D=d,S=o|X=x)}=$
$\left\{\frac{\Pr(Y=1,D=d,S=o\mid R=r,X=x)}{\Pr(Y=1,D=d,S=o|X=x)}-\frac{\Pr(Y=0,D=d,S=o\mid R=r,X=x)}{\Pr(Y=0,D=d,S=o|X=x)}\right\}\mu(d,x).$
@@
$\mathbb{E}\left[\frac{1\left\{D=d,S=e\right\}}{\Pr(D=d,S=e|X=x)}-\frac{1\left\{Y=0,D=d,S=o\right\}}{\Pr(Y=0,D=d,S=o|X=x)}\mid R=r,X=x\right]=$
$\mathbb{E}\left[\frac{1\left\{Y=1,D=d,S=o\right\}}{\Pr(Y=1,D=d,S=o|X=x)}-\frac{1\left\{Y=0,D=d,S=o\right\}}{\Pr(Y=0,D=d,S=o|X=x)}\mid R=r,X=x\right]\mu(d,x).$
@@ D.1 Algorithm details
    Learn the optimal representation on train: $\widehat{H}(R)$.
@@
      Learn the optimal representation: $\widehat{H}(R)=\frac{\widehat{\mathbb{E}}(\Delta^{o}|R)}{\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)}$ where $\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)$ is constructed from the marginal probabilities, predictors, and initial estimate via Lemma 4.1:
@@
    Construct an efficient causal estimate on test: $\widehat{\theta}$.
@@
      Construct an efficient causal estimate: $\widehat{\theta}=\frac{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{e}\widehat{H}(R)\}}{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{o}\widehat{H}(R)\}}$ where $\widehat{\Delta}^{e}$ and $\widehat{\Delta}^{o}$ are constructed from marginal probabilities according to (1):
      $\Delta^{e}=\frac{1\{D=1,S=e\}}{\textsc{count}_{D=1,S=e}}-\frac{1\{D=0,S=e\}}{\textsc{count}_{D=0,S=e}},\quad\Delta^{o}=\frac{1\{Y=1,S=o\}}{\textsc{count}_{Y=1,S=o}}-\frac{1\{Y=0,S=o\}}{\textsc{count}_{Y=0,S=o}}.$
@@
      Bootstrap its confidence interval: $\widehat{\theta}\pm c_{\alpha}\hat{v}n^{-1/2}$, where $c_{\alpha}$ is the $21-\alpha/$ 2 quantile of the standard Gaussian and $\hat{v}n^{-1/2}$ is the bootstrap standard error of $\widehat{\theta}$ while fixing $\widehat{H}(R)$.
D.2 Proof of Lemma 4.1
@@
$(\Delta^{e}-\Delta^{o}\theta)^{2}$ $=(\Delta^{e})^{2}-(2\theta)\Delta^{e}\Delta^{0}+\theta^{2}(\Delta^{o})^{2}=(\Delta^{e})^{2}+\theta^{2}(\Delta^{o})^{2}.$
@@
D.3 Proof of Proposition 4.1
To lighten notation, let $Y=\Delta^{e}$, $X=\Delta^{o}$, $U=Y-X\theta$, $Z=\widetilde{H}(R)$, and $\hat{Z}=\widehat{H}(R)$. With known marginal probabilities, the argument uses standard techniques, similar to Mackey et al., (2018); Chen et al., (2020). In this lighter notation,
@@
Focusing on the numerator, if $n^{1/2}\mathbb{E}_{n}(U\hat{Z})-n^{1/2}\mathbb{E}_{n}(UZ)=o_{p}(1)$ and $n^{1/2}\mathbb{E}_{n}(UZ)\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$, then $n^{1/2}\mathbb{E}_{n}(U\hat{Z})\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$ by Slutsky’s theorem .
Focusing on the denominator, if $\mathbb{E}_{n}(X\hat{Z})-\mathbb{E}_{n}(XZ)=o_{p}(1)$ and $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$ then $\mathbb{E}_{n}(X\hat{Z})=\mathbb{E}(XZ)+o_{p}(1)$ by the continuous mapping theorem.
@@
While proving the results, we use standard notation for cross fitting. Let there be $L$ folds, each denoted by $I_{\ell}$ with $\ell\in[L]$. Each fold contains $n_{\ell}=n/L$ observations. The complement of $I_{\ell}$ is $I_{-\ell}$. If $i\in I_{\ell}$, then $\hat{Z}_{i}=\widehat{H}_{\ell}(R_{i})$ is constructed from $\widehat{H}_{\ell}$ estimated on the remaining folds $I_{-\ell}$.
@@
Under Proposition 4.1’s conditions, $n^{1/2}\mathbb{E}_{n}(U\hat{Z})-n^{1/2}\mathbb{E}_{n}(UZ)=o_{p}(1)$.
@@
Due to cross fitting, the inner expectation is
@@
In the inequality, we use $E(U_{i}^{2}|Z_{i},I_{-\ell})=\mathbb{E}(U_{i}^{2}|Z_{i})\leq\bar{\sigma}^{2}_{U}$, where $\bar{\sigma}^{2}_{U}$ exists by hypothesis. In the final equality, we write $\mathcal{R}(\hat{Z})=\mathbb{E}\{(\hat{Z}_{i}-Z_{i})^{2}\mid I_{-\ell}\}=o_{p}(1)$ for the mean square limit. ∎
@@
Under Proposition 4.1’s conditions, $n^{1/2}\mathbb{E}_{n}(UZ)\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$.
@@
By Theorem 3.1, $\mathbb{E}(U_{i}Z_{i})=\mathbb{E}\{\mathbb{E}(U_{i}|Z_{i})Z_{i}\}=0$. Moreover, $\mathbb{E}(U_{i}^{2}Z_{i}^{2})=\mathbb{E}\{\mathbb{E}(U_{i}^{2}|Z_{i})Z_{i}^{2}\}\leq\bar{\sigma}^{2}_{U}\mathbb{E}(Z_{i}^{2})$ since $\mathbb{E}(U_{i}^{2}|Z_{i})\leq\bar{\sigma}^{2}_{U}$ under our assumptions. The latter is finite by hypothesis, so we apply the Lindeberg-Levy central limit theorem. ∎
@@
Under Proposition 4.1’s conditions, $\mathbb{E}_{n}(X\hat{Z})-\mathbb{E}_{n}(XZ)=o_{p}(1)$.
@@
We use $|X_{i}|\leq\bar{X}$ almost surely, which follows from our assumptions. ∎
@@
Under Proposition 4.1’s conditions, $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$.
@@
By Chebyshev’s inequality, it suffices to bound $\mathbb{V}(X_{i}Z_{i})\leq\mathbb{E}(X_{i}^{2}Z_{i}^{2})\leq\bar{X}^{2}\mathbb{E}(Z_{i}^{2})$. In summary, we use $|X_{i}|\leq\bar{X}$ almost surely and $\mathbb{E}(Z_{i}^{2})<\infty$. ∎
D.4 Proof of Proposition 4.2
With unknown marginal probabilities, some extra care is required. Extending the notation from the proof of Proposition 4.1, if $\hat{Y}=\widehat{\Delta}^{e}$, $\hat{X}=\widehat{\Delta}^{o}$, and $\hat{U}=\hat{Y}-\hat{X}\theta$, then

### v3

@@ B.3 Proof of Proposition 1
Appendix C Identification Proofs
C.1 Proof of Lemma 1
[…]
By the law of total probability,
$\delta_{d}^{e}(r,x)$ $=f_{R}(r\mid S=e,X=x,D=d)$
$=\int f_{R,Y}(r,y\mid S=e,X=x,D=d)\mathrm{d}y$
$=\int f_{R}(r\mid S=e,X=x,D=d,Y=y)f_{Y}(y\mid S=e,X=x,D=d)\mathrm{d}y.$
By Assumption 1, $f_{Y}(y\mid S=e,X=x,D=d)=f_{Y(d)}(y\mid S=e,X=x)$. Next, notice that
$f_{R}(r\mid S=e,X=x,D=d,Y=y)$ $=f_{R}(r\mid S=o,X=x,D=d,Y=y)$
$=f_{R}(r\mid S=o,X=x,Y=y)$
$=\delta_{y}^{o}(r,x),$
where the first equality applies Assumption 2 and the second equality applies Assumption 3(ii).
Combining the previous displays, we arrive at the general result
$\delta_{d}^{e}(r,x)=\int\delta_{y}^{o}(r,x)f_{Y(d)}(y\mid S=e,X=x)\mathrm{d}y.$ (3)
When $Y$ is binary,
$f_{Y(d)}(1\mid S=e,X=x)$ $=\mathbb{E}\{Y^{(d)}\mid S=e,X=x\}=\mu(d,x)$
$f_{Y(d)}(0\mid S=e,X=x)$ $=1-\mathbb{E}\{Y^{(d)}\mid S=e,X=x\}=1-\mu(d,x).$
Therefore, the general result specializes to
$\delta_{d}^{e}(r,x)=\delta_{1}^{o}(r,x)\mu(d,x)+\delta_{0}^{o}(r,x)\{1-\mu(d,x)\}=\delta_{0}^{o}(r,x)+\{\delta_{1}^{o}(r,x)-\delta_{0}^{o}(r,x)\}\mu(d,x).\quad\qed$
C.2 Proof of Theorem 1
Recall the notation from Lemma 1. To prove this result, we apply Bayes’ rule to rewrite
$\delta_{y}^{o}(r,x)$ $:=f_{R}(r\mid Y=y,S=o,X=x)=\frac{\Pr(Y=y,S=o\mid R=r,X=x)f_{R}(r|X=x)}{\Pr(Y=y,S=o|X=x)},$ (4)
$\delta_{d}^{e}(r,x)$ $:=f_{R}\left(r\mid D=d,S=e,X=x\right)=\frac{\Pr(D=d,S=e\mid R=r,X=x)f_{R}(r|X=x)}{\Pr(D=d,S=e|X=x)}.$
Applying Lemma 1 and canceling $f_{R}(r|X=x)$, we have for $d\in\{0,1\}$, $x\in\mathcal{X}$, and $r\in\mathcal{R}$,
$\frac{\Pr(D=d,S=e\mid R=r,X=x)}{\Pr(D=d,S=e|X=x)}-\frac{\Pr(Y=0,S=o\mid R=r,X=x)}{\Pr(Y=0,S=o|X=x)}$
$=\left\{\frac{\Pr(Y=1,S=o\mid R=r,X=x)}{\Pr(Y=1,S=o|X=x)}-\frac{\Pr(Y=0,S=o\mid R=r,X=x)}{\Pr(Y=0,S=o|X=x)}\right\}\mu(d,x).$
@@ C.2 Proof of Theorem 1
$\mathbb{E}\left[\frac{1\left\{D=d,S=e\right\}}{\Pr(D=d,S=e|X=x)}-\frac{1\left\{Y=0,S=o\right\}}{\Pr(Y=0,S=o|X=x)}\mid R=r,X=x\right]$
$=\mathbb{E}\left[\frac{1\left\{Y=1,S=o\right\}}{\Pr(Y=1,S=o|X=x)}-\frac{1\left\{Y=0,S=o\right\}}{\Pr(Y=1,S=o|X=x)}\mid R=r,X=x\right]\mu(d,x).$
@@
$\mathbb{E}\left[\frac{1\left\{D=1,S=e\right\}}{\Pr(D=1,S=e|X=x)}-\frac{1\left\{D=0,S=e\right\}}{\Pr(D=0,S=e|X=x)}\mid R=r,X=x\right]$
$=\mathbb{E}\left[\frac{1\left\{Y=1,S=o\right\}}{\Pr(Y=1,S=o|X=x)}-\frac{1\left\{Y=0,S=o\right\}}{\Pr(Y=1,S=o|X=x)}\mid R=r,X=x\right]\theta(x),$
@@
C.3 Proof of Theorem 2
The proof follows the same steps as the proofs of Lemma 1 and Theorem 1.
  1. 1.
    As in Lemma 1, for general $(X,Y,R)$ and for $\delta_{d}^{e}(r,x):=f_{R}(r\mid S=e,X=x,D=d)$,
    $\delta_{d}^{e}(r,x)=\int f_{R}(r\mid S=e,X=x,D=d,Y=y)f_{Y}(y\mid S=e,X=x,D=d)\mathrm{d}y.$
    By Assumption 1, $f_{Y}(y\mid S=e,X=x,D=d)=f_{Y(d)}(y\mid S=e,X=x)$. Now, however,
    $f_{R}(r\mid S=e,X=x,D=d,Y=y)$ $=f_{R}(r\mid S=o,X=x,D=d,Y=y)=:\delta_{y,d}^{o}(r,x),$
    where the first equality applies Assumption 2 and the second equality applies Assumption 3(i).
    Combining the previous displays, we arrive at the general result
    $\delta_{d}^{e}(r,x)=\int\delta_{y,d}^{o}(r,x)f_{Y(d)}(y\mid S=e,X=x)\mathrm{d}y.$ (5)
    When $Y$ is binary, as before
    $f_{Y(d)}(1\mid S=e,X=x)$ $=\mu(d,x),\quad f_{Y(d)}(0\mid S=e,X=x)=1-\mu(d,x).$
    Therefore, the general result specializes to
    $\delta_{d}^{e}(r,x)=\delta_{1,d}^{o}(r,x)\mu(d,x)+\delta_{0,d}^{o}(r,x)\{1-\mu(d,x)\}=\delta_{0,d}^{o}(r,x)+\mu(d,x)\{\delta_{1,d}^{o}(r,x)-\delta_{0,d}^{o}(r,x)\}.$
  2. 2.
    As in Theorem 1, we next apply Bayes’ rule to rewrite
    $\delta_{y,d}^{o}(r,x)$ $=\frac{\Pr(Y=y,D=d,S=o\mid R=r,X=x)f_{R}(r|X=x)}{\Pr(Y=y,D=d,S=o|X=x)}$ (6)
    $\delta_{d}^{e}(r,x)$ $=\frac{\Pr(D=d,S=e\mid R=r,X=x)f_{R}(r|X=x)}{\Pr(D=d,S=e|X=x)}.$
    It therefore follows by canceling $f_{R}(r|X=x)$ that, for $d\in\{0,1\}$ and $r\in\mathcal{R}$,
    $\frac{\Pr(D=d,S=e\mid R=r,X=x)}{\Pr(D=d,S=e|X=x)}-\frac{\Pr(Y=0,D=d,S=o\mid R=r,X=x)}{\Pr(Y=0,D=d,S=o|X=x)}$
    $=\left\{\frac{\Pr(Y=1,D=d,S=o\mid R=r,X=x)}{\Pr(Y=1,D=d,S=o|X=x)}-\frac{\Pr(Y=0,D=d,S=o\mid R=r,X=x)}{\Pr(Y=0,D=d,S=o|X=x)}\right\}\mu(d,x).$
@@ C.3 Proof of Theorem 2
    $\mathbb{E}\left[\frac{1\left\{D=d,S=e\right\}}{\Pr(D=d,S=e|X=x)}-\frac{1\left\{Y=0,D=d,S=o\right\}}{\Pr(Y=0,D=d,S=o|X=x)}\mid R=r,X=x\right]$
    $=\mathbb{E}\left[\frac{1\left\{Y=1,D=d,S=o\right\}}{\Pr(Y=1,D=d,S=o|X=x)}-\frac{1\left\{Y=0,D=d,S=o\right\}}{\Pr(Y=0,D=d,S=o|X=x)}\mid R=r,X=x\right]\mu(d,x).$
@@ D.1 Algorithm Details
    Learn the representation on train: $\widehat{H}(R)$.
@@
      Learn the representation: $\widehat{H}(R)=\frac{\widehat{\mathbb{E}}(\Delta^{o}|R)}{\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)}$ where $\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)$ is constructed from the marginal probabilities, predictors, and initial estimate via Lemma 2:
@@
    Construct a causal estimate on test: $\widehat{\theta}$.
@@
      Construct a causal estimate: $\widehat{\theta}=\frac{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{e}\widehat{H}(R)\}}{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{o}\widehat{H}(R)\}}$ where $\widehat{\Delta}^{e}$ and $\widehat{\Delta}^{o}$ are constructed from marginal probabilities according to (1):
      $\widehat{\Delta}^{e}=\frac{1\{D=1,S=e\}}{\textsc{count}_{D=1,S=e}}-\frac{1\{D=0,S=e\}}{\textsc{count}_{D=0,S=e}},\quad\widehat{\Delta}^{o}=\frac{1\{Y=1,S=o\}}{\textsc{count}_{Y=1,S=o}}-\frac{1\{Y=0,S=o\}}{\textsc{count}_{Y=0,S=o}}.$
@@
      Bootstrap its confidence interval: $\widehat{\theta}\pm c_{\alpha}\hat{v}n^{-1/2}$, where $c_{\alpha}$ is the $1-\alpha/2$ quantile of the standard Gaussian and $\hat{v}n^{-1/2}$ is the bootstrap standard error of $\widehat{\theta}$ while fixing $\widehat{H}(R)$.
D.2 Proof of Lemma 2
@@
$(\Delta^{e}-\Delta^{o}\theta)^{2}$ $=(\Delta^{e})^{2}-(2\theta)\Delta^{e}\Delta^{o}+\theta^{2}(\Delta^{o})^{2}=(\Delta^{e})^{2}+\theta^{2}(\Delta^{o})^{2}.$
@@
D.3 Proof of Proposition 2
To lighten notation, let $Y=\Delta^{e}$, $X=\Delta^{o}$, $U=Y-X\theta$, $Z=\widetilde{H}(R)$, and $\hat{Z}=\widehat{H}(R)$. By construction, $\mathbb{E}(U^{2}|Z)\leq\bar{\sigma}^{2}_{U}$ for some constant $\bar{\sigma}^{2}_{U}<\infty$ due to Lemma 2 and the assumption that marginal probabilities are bounded away from zero. Moreover, $|X|\leq\bar{X}$ almost surely for some constant $\bar{X}<\infty$ due to the assumption that marginal probabilities are bounded away from zero.
With known marginal probabilities, the argument uses standard techniques, similar to Mackey et al. (2018); Chen et al. (2020). In this lighter notation,
@@ D.3 Proof of Proposition 2
Focusing on the numerator, if $n^{1/2}\mathbb{E}_{n}(U\hat{Z})=n^{1/2}\mathbb{E}_{n}(UZ)+o_{p}(1)$ and $n^{1/2}\mathbb{E}_{n}(UZ)\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$, then $n^{1/2}\mathbb{E}_{n}(U\hat{Z})\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$ by Slutsky’s theorem .
Focusing on the denominator, if $\mathbb{E}_{n}(X\hat{Z})=\mathbb{E}_{n}(XZ)+o_{p}(1)$ and $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$ then $\mathbb{E}_{n}(X\hat{Z})=\mathbb{E}(XZ)+o_{p}(1)$ by the continuous mapping theorem.
@@
While proving the results, we use standard notation for cross-fitting. Let there be $L$ folds, each denoted by $I_{\ell}$ with $\ell\in[L]$. Each fold contains $n_{\ell}=n/L$ observations. The complement of $I_{\ell}$ is $I_{-\ell}$. If $i\in I_{\ell}$, then $\hat{Z}_{i}=\widehat{H}_{\ell}(R_{i})$ is constructed from $\widehat{H}_{\ell}$ estimated on the remaining folds $I_{-\ell}$.
@@
Under Proposition 2’s conditions, $n^{1/2}\mathbb{E}_{n}(U\hat{Z})=n^{1/2}\mathbb{E}_{n}(UZ)+o_{p}(1)$.
@@
Due to cross-fitting, the inner expectation is
@@
In the inequality, we use $\mathbb{E}(U_{i}^{2}|Z_{i},I_{-\ell})=\mathbb{E}(U_{i}^{2}|Z_{i})\leq\bar{\sigma}^{2}_{U}$, where $\bar{\sigma}^{2}_{U}<\infty$ under our assumptions. In the final equality, we write $\mathcal{R}(\hat{Z})=\mathbb{E}\{(\hat{Z}_{i}-Z_{i})^{2}\mid I_{-\ell}\}=o_{p}(1)$ for the mean square limit. ∎
@@
Under Proposition 2’s conditions, $n^{1/2}\mathbb{E}_{n}(UZ)\rightsquigarrow\mathcal{N}\{0,\mathbb{E}(U^{2}Z^{2})\}$.
@@
By Theorem 1, $\mathbb{E}(U_{i}Z_{i})=\mathbb{E}\{\mathbb{E}(U_{i}|Z_{i})Z_{i}\}=0$. Moreover, $\mathbb{E}(U_{i}^{2}Z_{i}^{2})=\mathbb{E}\{\mathbb{E}(U_{i}^{2}|Z_{i})Z_{i}^{2}\}\leq\bar{\sigma}^{2}_{U}\mathbb{E}(Z_{i}^{2})$ since $\mathbb{E}(U_{i}^{2}|Z_{i})\leq\bar{\sigma}^{2}_{U}$ under our assumptions. Here, $\mathbb{E}(Z_{i}^{2})$ is finite by Assumption 4, so we apply the Lindeberg-Levy central limit theorem. ∎
@@
Under Proposition 2’s conditions, $\mathbb{E}_{n}(X\hat{Z})=\mathbb{E}_{n}(XZ)+o_{p}(1)$.
@@
We use $|X_{i}|\leq\bar{X}$ almost surely, since marginal probabilities are bounded away from zero. ∎
@@
Under Proposition 2’s conditions, $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$.
@@
By Chebyshev’s inequality, it suffices to bound $\mathbb{V}(X_{i}Z_{i})\leq\mathbb{E}(X_{i}^{2}Z_{i}^{2})\leq\bar{X}^{2}\mathbb{E}(Z_{i}^{2})$. In summary, we use $|X_{i}|\leq\bar{X}$ almost surely since marginal probabilities are bounded away from zero and $\mathbb{E}(Z_{i}^{2})<\infty$ due to Assumption 4. ∎
D.4 Proof of Proposition 3
With unknown marginal probabilities, some extra care is required. Extending the notation from the proof of Proposition 2, if $\hat{Y}=\widehat{\Delta}^{e}$, $\hat{X}=\widehat{\Delta}^{o}$, and $\hat{U}=\hat{Y}-\hat{X}\theta$, then